OpenStudy (anonymous):

At 700 K, a 2.00-L flask contains 0.00934 mol of H2 , 0.00516 mol of I2 , and 0.051 mol of HI at equilibrium. What is the equilibrium constant of the given reaction at this temperature?

OpenStudy (matt101):

To calculate the equilibrium constant, you'll need the balanced chemical reaction. I'm assuming the reaction you want is this: \[H_2 + I_2 \rightarrow 2HI\] Remember to divide each mole value by the volume to get the concentration. Then use those values in the K(eq) equation: \[K_{eq}={[HI]^2 \over [H_2][I_2]}\] \[K_{eq}={{\left( 0.051 \over 2 \right)^2} \over {\left( 0.00934 \over 2 \right) \left(0.00516 \over 2 \right)}}\] \[K_{eq} \approx 54\]